OpenStudy (anonymous):

Find the variance and standard deviation for the following set of data: 101,102,100,100,110,109,109,108,109 I've tried a hundred different ways to solve this and I can never get an accurate answer. Can someone please explain this to me? Thanks (:

OpenStudy (amistre64):

we can make this less daunting by moving the range of values to more workable values subtract 100 from each value and it works out the same

OpenStudy (amistre64):

\[X:\{x_1+k,x_2+k,x_3+k,...,x_n+k\}\] \[mean(X):\frac{x_1+k+x_2+k+x_3+k+...+x_n+k}{n}\] \[mean(X):\frac{(x_1+x_2+x_3+...+x_n)+kn}{n}\] \[mean(X):\frac{(x_1+x_2+x_3+...+x_n)}{n}+k\] 1,2,0,0,10,9,9,8,9 what is the mean of this equivalent set?

OpenStudy (amistre64):

moving the range of the data to something more workable doesnt change the variance, and since variance is the square of standard deviation, it doesnt alter the sd either can you work the var and sd with this set now?

OpenStudy (anonymous):

Yes thank you (:

OpenStudy (amistre64):

cool :) if you still need help, feel free to ask